170829 UCSC

The fog collector came on at 12:48 am last night and turned off at 10:53 am this morning. 180mL of sample was collected. The collector was washed with soap and water and a blank was taken. Brian set up another light incubation experiment. Hopefully, non of the jars break this time!

Leave a Reply

Your email address will not be published. Required fields are marked *